OpenStudy (anonymous):

Use the five-step strategy for solving word problems to find the number or numbers described in the following exercise. When a number is decreased by 25% of itself, the result is 216. What is the number? A) 72 B) 1152 C) 288 D) 9

OpenStudy (anonymous):

x-25%=216

OpenStudy (anonymous):

x-25%x=216 (but 25% is written as .25)

OpenStudy (anonymous):

k